OpenStudy (anonymous):

simplify the following: (c - 2) (c - 5)

OpenStudy (anonymous):

c^2 -7c + 10

OpenStudy (anonymous):

c^2-5c-3c+10 c^2-8c+10

OpenStudy (anonymous):

Remember the FOIL equation. First, Outer, Inner, Last

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

min{(c-2) (c-5)} = -9/4 at c = 7/2